OpenStudy (anonymous):

A train crosses a pole in 15 seconds, and it crosses a 100 meters platform in 25 seconds, then what is the Length of the train.

OpenStudy (anonymous):

\[\frac x {15} =\frac {x+100}{25} \implies x = 150\]

OpenStudy (anonymous):

no ... at first even I did the same ... But the answer comes out to be 159m.

OpenStudy (anonymous):

The right answer is 150 m.

OpenStudy (apoorvk):

Yeah the right answer should be 150 metres. There is a printing error in your question source it seems

OpenStudy (anonymous):

@lgbasallote What's ur opinion

OpenStudy (lgbasallote):

well according to @FoolForMath equation it should be 150

OpenStudy (anonymous):

@ParthKohli need some help here...

Parth (parthkohli):

FoolForMath is correct. \(\Large \color{purple}{\rightarrow 25x = 15( x + 100) }\) \(\Large \color{purple}{\rightarrow 25x = 15x + 1500 }\) \(\Large \color{purple}{\rightarrow 10x = 1500 }\) \(\Large \color{purple}{\rightarrow x = 150 }\) Maybe, you had committed a mistake, if you got 159

OpenStudy (anonymous):

no the answer given is 159m.

Parth (parthkohli):

I don't see any mistake in @FoolForMath's answer. Maybe it's a misprint as 9 is close to 0 on the keyboard.

OpenStudy (apoorvk):

Let's just explain what's happening in here. In case of 'crossing' the pole the train is basically just travelling its own length. Let its own length be 'x'. So speed of train v = x/(time taken) = x/15 ...{i} When the train is crossing the platform, it's travelling both its own length plus the length of the platform. so it travels (x+100) metres. here, v= (x+100)/25 ....{ii} From 'i' and 'ii', you can only arrive at x=150 the answer is definitely wrong - it's quite probable that the person typing it out pressed a '9' instead of '0'. Lok at their positions on your keyboard :)

OpenStudy (anonymous):

May be thanks...

OpenStudy (anonymous):

Oh come on @lgbasallote you were replying something

Parth (parthkohli):

 After the engine crosses the end of the platform, then the rest of the train must cross the length of the train to make the complete train pass through.